OpenStudy (butterflyhope):

Urgent (Medal): Write algebraic equations expressing the two identity properties for the set of real numbers. Thanks :)

OpenStudy (anonymous):

Additive identity property: a+0=0+a =a Multiplicative identity property: 1.a=a.1=a